# CATArena: Evaluation of LLM Agents Through Iterative Tournament Competitions

### v1

@@ 5 Conclusion
In this work, we address two fundamental challenges in LLM agent evaluation: the need for systematic measurement of learning ability, and the tendency of traditional benchmarks to become saturated as agent capabilities improve. To this end, we propose an iterative peer-learning-based competitive framework, enabling agents to continually revise and enhance their strategies through dynamic interaction and feedback. Building on this, we introduce CATArena, a tournament-style benchmark featuring open-ended and rankable board and card games. CATArena provides an environment with unlimited potential for agent improvement potential and extensible evaluation across new domains. Experimental results demonstrate that our framework reliably assesses core agent abilities, particularly learning ability and strategy coding, while ensuring stability and scalability. The open and flexible architecture of CATArena supports ongoing research and benchmarking for future intelligent agents.
Limitations.
The current evaluation in CATArena is limited to four games, which primarily assess agents’ learning ability and strategy coding. These scenarios do not encompass the full spectrum of potential LLM agent capabilities. In future work, we plan to introduce a wider variety of more complex tasks to evaluate agents’ learning and other abilities from different perspectives.
Appendix A Tournament Format and Scoring System
Table 6: Configs of tournament on four games.
Environment Code Agents LLM-player
Gomoku
Board size: $\times15$
Number of pairwise matches: 4 $\times$ 2
Swap black and white pieces after each match
Maximum time per move: 10 s
Board size: $\times15$
Number of pairwise matches: 2 $\times$ 2
Swap black and white pieces after each match
Maximum time per move: 600 s
Texas Hold’em
Max players: 12
Rounds: 100
Random shuffle seat after each round
Initial chips: 2000
Blind increase every 24 hands
Max hands per round: 720 or until winner decided
Maximum time per move: 3 s
Max players: 12
Rounds: 10
Random shuffle seat after each round
Initial chips: 2000
Blind increase every 24 hands
Max hands per round: 720 or until winner decided
Maximum time per move: 1000 s
Bridge
Number of pairwise matches: 12 $\times$ 2
Swap directions of open/closed rooms
Use same deck for each pair of match
Maximum time per move: 10 s
Number of pairwise matches: 12 $\times$ 2
Swap directions of open/closed rooms
Use same deck for each pair of match
Maximum time per move: 200 s
Chess
Number of pairwise matches: 8 $\times$ 2
Swap black and white pieces after each match
Maximum moves per game: 200
Maximum time per move: 10 s
Number of pairwise matches: 2 $\times$ 2
Swap black and white pieces after each match
Maximum moves per game: 200
Maximum time per move: 600 s
We list the basic settings for each games in Table 6.
For each game, we ensure that the number of pairwise matches among code agents allows the final results to stabilize, i.e., for each game, the L1-norm fluctuation of the scoring matrix $W$ is less than 5%.
For LLM-players, since their reasoning time is relatively long for most games, we reduce the number of repeated experiments. Note that our comparison with LLM-players is only a qualitative analysis of the differences between LLM-player and corresponding coding agent. The exploration of LLM-players’ results is not the focus of this paper; refer to the main text for details.
Appendix B Evaluation Metric Calculation
Table 7: Scoring rules of tournament on four games.
@@ Appendix B Evaluation Metric Calculation
To evaluate the basic capabilities of LLM agents, we define metrics for each applied game. Let $N$ be the number of rounds and $K$ be the number of participating agents. We construct a matchup matrix
$W\in\mathbb{R}^{(N\cdot K)\times(N\cdot K)},$
@@
For asymmetric games, pairwise results are not feasible; instead, batch-based tournaments are used and the score matrix records the win rates of multi-agent matches. For each batch, we obtain a single score group $W_{i_{1},i_{2},\ldots,i_{BS}}^{n_{1},n_{2},\ldots,n_{BS}}$, where $BS$ is the batch size. We conduct three types of experiments to accommodate different metric calculations: (1) $W_{i}^{1,2,\ldots,N}$, where the same agent’s strategies from $N$ rounds compete against each other, used to compute self-improvement metrics $S_{i}^{n}$; (2) $W_{i_{1},i_{2},\ldots,i_{T}}^{n}$, where all agents in the same round compete, used to calculate the base score $B$; (3) All $N\times T$ agent strategies are randomly shuffled and grouped for competition (with $BS=12$ in our experiments), used to compute both the global score $G_{i}^{n}$ and advanced score $A_{i}^{n}$.
Scoring rules for the four games are summarized in Table 7.
Appendix C Generation Configs
@@
For all LLMs used in our work, we set temperature to be $0.1$, max token identical to their official APIs’ setting. We set top-p to $1.0$, Top-k to be $100$, and presence penalty to be default to the API.
@@
For LLM-players, considering the uncertainty in model output formats, we allow up to three retries. The prompt of LLM-players are in arena’s code and not present in paper considering its excessive length.
Appendix D Repetition Experiments
@@
Table 8: Standard deviation of ranking in Round 1 and Round 2 with repeating 4 times.
Games Average Gomoku Hold’em Bridge Chess
Standard Variant Standard Variant Standard Variant Standard Variant Standard Variant
Round 1
Minimal
Claude-4-Sonnet 0.80 0.91 1.58 0.43 1.12 0.71 0.50 1.64 0.00 0.87
Deepseek-Chat 0.72 0.81 0.83 0.87 0.83 1.22 1.22 0.71 0.00 0.43
Doubao-Seed 1.58 0.90 1.87 1.87 1.09 0.43 2.06 0.87 1.30 0.43
Gemini-2.5-Pro 1.24 1.23 1.50 1.30 1.12 0.83 1.12 1.66 1.22 1.12
GPT-5 0.75 1.18 0.71 1.50 0.71 1.30 1.09 1.50 0.50 0.43
Qwen3-Coder 1.16 0.84 1.48 1.50 1.22 0.71 1.09 0.71 0.83 0.43
Commercial
Claude-Code 1.27 1.01 1.12 1.66 0.43 0.00 1.79 1.09 1.73 1.30
CodeX 0.76 0.57 0.83 0.50 0.43 0.50 1.09 1.30 0.71 0.00
Gemini-CLI 1.14 0.93 0.83 0.43 1.12 1.00 1.30 1.79 1.30 0.50
Qwen-Coder 1.01 0.95 1.22 1.48 1.12 0.00 0.87 0.83 0.83 1.50
Round 2
Minimal
Claude-4-Sonnet 0.81 0.55 1.48 0.50 0.83 0.43 0.50 0.83 0.43 0.43
Deepseek-Chat 0.94 1.03 1.30 1.09 0.87 0.83 1.09 1.48 0.50 0.71
Doubao-Seed 0.79 0.88 1.09 1.12 0.87 0.71 0.71 0.87 0.50 0.83
Gemini-2.5-Pro 1.30 1.28 1.66 1.92 1.00 1.22 2.06 1.12 0.50 0.87
GPT-5 0.89 1.22 1.00 1.58 0.43 1.66 1.64 1.22 0.50 0.43
Qwen3-Coder 0.82 1.02 1.58 1.30 0.83 0.43 0.43 1.50 0.43 0.83
Commercial
Claude-Code 1.20 1.04 1.09 1.41 1.22 0.43 1.64 1.09 0.83 1.22
CodeX 0.71 0.87 0.43 0.87 0.50 1.09 0.83 1.09 1.09 0.43
Gemini-CLI 0.88 1.10 0.83 1.64 0.83 0.83 1.41 1.48 0.43 0.43
Qwen-Coder 1.13 0.90 1.66 0.83 0.43 0.83 1.00 1.12 1.41 0.83
We report the results of repetition experiments($N=4$) on first two tournament round in Table 8.
From table, we observe that 1. The rankings of most agents remain relatively stable ,with ranking standard deviation changes of less than one. However, a few agents, such as Gemini-2.5-Pro and Claude-Code, exhibit greater fluctuations; 2. The rankings for standard games are more stable than those for variant games; 3. The results of the open source model are more stable than those of the closed source model, and commercial agents are more stable than minimal agents; 4. Additionally, we observe that agents do not consistently generate runnable code repositories across multiple development attempts. Even commercial agents occasionally fail to produce successful builds, which suggests that current code agents still need to improve their development stability.
Appendix E Learning Ability
E.1 Global Learning Trend

### v2

@@ 5 Conclusion
In this paper, we present CATArena, a comprehensive framework that shifts the evaluation of code agents from static correctness to dynamic evolutionary capability. By simulating the iterative nature of real-world software engineering within a competitive tournament structure, CATArena quantifies the ability of agents to leverage execution feedback and peer knowledge for continuous optimization. Through extensive evaluations on CATArena, we benchmark the evolutionary capabilities of state-of-the-art LLM code agents, confirming that evolutionary potential is a distinct dimension from single-turn generation capability. Our in-depth analysis further dissects these evolutionary patterns, revealing the specific contributions of self-reflection and peer-learning mechanisms to agent growth. Moreover, we demonstrate that evolutionary trajectories are highly task-dependent; CATArena’s diverse spectrum of competitive and objective tasks proves essential for capturing these nuanced behavioral shifts. In the future, CATArena will integrate more tasks and serve as a highly extensible testbed.
Impact Statement
This paper presents CATArena, a benchmark for evaluating the evolutionary capabilities of code agents. Our goal is to advance the reliability and efficiency of automated software engineering. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Evaluation Metric Calculation
Table 5: Scoring rules of $W$ matrix calculation on six tasks.
@@ Appendix A Evaluation Metric Calculation
OIBench TimeAUC per problem: CDF of $\log_{10}(t/T_{\mathrm{ref}})$ AUC over $x_{\min},1]$, normalized by reference AUC.
SWE-Perf Performance: mean significant speedup across efficiency tests (Mann-Whitney U, p<0.1p{<}0.1) vs baseline.
To provide a unified evaluation standard across diverse tasks, we construct a global performance matrix $W$. Scoring rules for the tasks are summarized in Table 5. Let $N$ be the number of rounds and $M$ be the number of participating agents. We construct a matchup matrix
$W\in\mathbb{R}^{(N\cdot M)\times(N\cdot M)},$
@@
For Hold’em, pairwise results are not feasible; instead, batch-based tournaments are used and the score matrix records the win rates of multi-agent matches. For each batch, we obtain a single score group $W_{i_{1},i_{2},\ldots,i_{BS}}^{n_{1},n_{2},\ldots,n_{BS}}$, where $BS$ is the batch size. We conduct three types of experiments to accommodate different metric calculations: (1) $W_{i}^{1,2,\ldots,N}$, where the same agent’s strategies from $N$ rounds compete against each other, used to compute self-improvement metrics $S_{i}^{n}$; (2) $W_{i_{1},i_{2},\ldots,i_{T}}^{n}$, where all agents in the same round compete, used to calculate the base score $B$; (3) All $N\times T$ agent strategies are randomly shuffled and grouped for competition (with $BS=12$ in our experiments), used to compute both the global score $G_{i}^{n}$ and advanced score $A_{i}^{n}$.
For objective tasks, we simplify the notation $W_{i,j}^{n,m}$ to $W_{i}^{n}$.
Appendix B Tournament Format and Repetition
Table 6: Configs of tournament on six tasks.
Task Type Environment Code Agents
Competitive Gomoku
Board size: $\times15$
Number of pairwise matches: 4 $\times$ 2
Swap black and white pieces after each match
Maximum time per move: 10 s
Competitive Texas Hold’em
Max players: 12
Rounds: 100
Randomly shuffle seats after each round
Initial chips: 2000
Blind increase every 24 hands
Max hands per round: 720 or until winner decided
Maximum time per move: 3 s
Competitive Bridge
Number of pairwise matches: 12 $\times$ 2
Swap directions of open/closed rooms
Use same deck for each pair of matches
Maximum time per move: 10 s
Competitive Chess
Number of pairwise matches: 8 $\times$ 2
Swap black and white pieces after each match
Maximum moves per game: 200
Maximum time per move: 10 s
Objective OIBench
Problems per round: 5
Language: C++
Per-test timeout: 2s
Score: mean TimeAUC
Agent time budget: 60 min
Objective SWE-Perf
Instance: repo@base_commit (5 instances per sweep)
Efficiency tests from dataset
[…]
Score: mean significant speedup over baseline
Figure 5: Cumulative win rate trajectories of Hold’em agents over 300 tournament rounds.
We list the basic settings for each task in Table 6.
For competitive tasks, we ensure that the number of pairwise matches among code agents allows the final results to stabilize, i.e., for each task, the L1-norm fluctuation of the scoring matrix $W$ is less than 5%.
For objective tasks, agents are evaluated independently on the same fixed benchmark each round (no pairwise matches). Each agent submits code, the arena runs the official evaluator to compute $W_{i}^{n}$, and a round leaderboard is produced. The leaderboard, logs, and top solutions/patches are fed back to all agents for the next round, forming an iterative optimization tournament over $N$ rounds with a diagonal score matrix.
Specifically, due to the inherent randomness in card dealing in Texas Hold’em, we conduct an additional analysis on the stability of the win rates. Figure 5 illustrates the cumulative win rate trends over 10 to 300 rounds for a representative tournament. To further assess the reliability of our evaluation, we analyze the win rate trends from different starting rounds to the final round. Based on this analysis, starting from round 50, the cumulative win rates of all models become stable, with the absolute value of the linear regression slope dropping below 0.0002 (i.e., per-round changes less than 0.02%). This indicates that our tournament design of 100 rounds, with each round consisting of 200 to 720 hands, is sufficient to cover the stochasticity of the game and provides a robust and reliable assessment of model performance. The results of the tournament reliably reflect the true capabilities of the evaluated models, ensuring the soundness of subsequent algorithm comparisons and optimizations.
Appendix C Comparative Analysis of Elo Rating and Global Win Rate
(a) Elo Rating History
(b) Win Rate History
Figure 6: Comparison of performance metrics in a 100-game Texas Hold’em tournament among 6 random selected players. (a) The Elo rating fluctuates significantly due to match order sensitivity. (b) The global win rate converges to a stable value, providing a more reliable foundation for calculating the evolutionary growth rate.
In this section, we justify our choice of the global average win rate (the basis for our $G$ score) over the Elo rating system by comparing their stability in a 100-game Texas Hold’em tournament. As visualized in Figure 6, the Elo rating (a) exhibits significant volatility and sensitivity to match order, characterized by "zigzag" oscillations that reflect transient outcomes rather than intrinsic capability. This "recency bias" makes Elo unsuitable for capturing a stable performance snapshot for each evolutionary round. In contrast, the Global Win Rate (b) demonstrates smooth convergence and effectively filters out stochastic noise. By reflecting the agent’s expected performance against the entire population, the win rate provides a robust and consistent metric, ensuring that the calculated evolutionary slope represents genuine strategic growth rather than artifacts of tournament scheduling.
Appendix D Generation Configs
Table 7: Agent Specifications and Open-Source Status.
Agent Type Agent Framework Model Agent OSS LLM OSS
Qwen3-Coder-480B (Qwen, 2025b)
Doubao-Seed-1.6 (Seed, 2025)
GPT-5 (OpenAI, 2025b)
Claude-4-Sonnet (Anthropic, 2025b)
Gemini-2.5-Pro (Gemini, 2025)
Commercial Gemini-CLI (Google, 2025) Gemini-2.5-Pro (Gemini, 2025)
Claude-Code (Anthropic, 2025a) Claude-4/3.7 Hybrid (Anthropic, 2025b)
Codex (OpenAI, 2025a) GPT-5 (OpenAI, 2025b)
Qwen-Code (Qwen, 2025a) Qwen3-Coder-480B (Qwen, 2025b)
@@ Appendix D Generation Configs
For all LLMs used in our work, we set the temperature to 0.1 and the max tokens to the official API defaults. We set top-p to 1.0, top-k to 100, and the presence penalty to the API default.
@@
Appendix E Repetition Experiments: Stability Analysis
@@
Table 8: Standard deviation of agent ranking in Round 1 and Round 2 with repeating 4 times.
Games Gomoku Hold’em Bridge Chess OIBench SWE-Perf
Round 1 Minimal Claude-4-Sonnet 1.58 1.12 0.50 0.00 0.50 0.82
Deepseek-Chat 0.83 0.83 1.22 0.00 0.50 1.83
Doubao-Seed 1.87 1.09 2.06 1.30 0.96 2.00
Gemini-2.5-Pro 1.50 1.12 1.12 1.22 0.58 0.96
GPT-5 0.71 0.71 1.09 0.50 0.00 0.50
Qwen3-Coder 1.48 1.22 1.09 0.83 0.50 0.82
Comm. Claude-Code 1.12 0.43 1.79 1.73 0.96 -
Codex 0.83 0.43 1.09 0.71 1.41 -
Gemini-CLI 0.83 1.12 1.30 1.30 0.82 -
Qwen-Coder 1.22 1.12 0.87 0.83 0.96 -
Round 2 Minimal Claude-4-Sonnet 1.48 0.83 0.50 0.43 0.50 1.23
Deepseek-Chat 1.30 0.87 1.09 0.50 0.58 0.00
Doubao-Seed 1.09 0.87 0.71 0.50 2.16 0.00
Gemini-2.5-Pro 1.66 1.00 2.06 0.50 0.82 1.50
GPT-5 1.00 0.43 1.64 0.50 1.00 1.16
Qwen3-Coder 1.58 0.83 0.43 0.43 0.50 0.00
Comm. Claude-Code 1.09 1.22 1.64 0.83 0.58 -
Codex 0.43 0.50 0.83 1.09 1.91 -
Gemini-CLI 0.83 0.83 1.41 0.43 1.73 -
Qwen-Coder 1.66 0.43 1.00 1.41 0.58 -
Table 8 presents the results of four independent repetitions of the first two tournament rounds. The primary objective is to assess whether stochasticity in code generation significantly alters the competitive landscape.
The data reveals that ranking stability is robust: the majority of agents maintain their relative positions with a standard deviation of less than one rank. While specific implementations vary per run, the performance gap between agents is sufficiently large to maintain a stable leaderboard.
Other minor observations include: (1) standard games exhibit higher stability than variant games; (2) open-source and commercial agents generally outperform minimal agents in terms of consistency. Finally, we observe that "runnability" is not guaranteed across runs, suggesting that while the relative capability (ranking) is stable, the absolute success rate of code generation still suffers from inherent randomness.
Appendix F Error Rate Analysis
Table 9: Proportion of HTTP errors encountered during code execution for each agent in Round 1.
Gomoku Hold’em Bridge Chess
HTTP error(%) Standard Variant Standard Variant Standard Variant Standard Variant
Minimal Claude-4-Sonnet 7 0 0 0 0 0 0 0
Deepseek-Chat 21 0 0 48 0 0 100 100
Doubao-Seed 17 0 49 93 12 0 0 100
Gemini-2.5-Pro 3 100 8 50 0 1 0 0
GPT-5 6 0 0 0 0 100 2 21
Qwen3-Coder 11 0 0 91 2 1 0 0
Commercial best ADK 0 0 0 0 0 0 0 0
Claude-Code 0 0 0 0 0 0 0 0
Codex 0 0 0 0 0 0 0 0
Gemini-CLI 0 0 0 0 48 0 0 0
Qwen-Coder 2 0 0 18 0 17 0 0
During the implementation of strategy code, some agents encounter runtime errors. We observe that all agents were able to generate the required files successfully. However, most errors occur in interface implementation and action validity. These cases were consistently recorded as HTTP errors. Whenever an HTTP error occurs, a default action is automatically executed to ensure the continuity of the competition. For Texas Hold’em Poker, the default action is to fold for that hand. For Chess and Gomoku, the default action is for the player to concede the current game.
Table 9 reports the proportion of HTTP error actions encountered during the code development of each agent in Round 1. The error rate is calculated as the rate of failed actions in total number of actions taken by the strategy code. The experimental results indicate that commercial agents exhibit a significantly lower failure rate compared to minimal agents. Additionally, the failure rate is higher in variant games than in traditional games.
Appendix G Detailed Evolutionary Trajectory Analysis
Figure 7: Trend of global performance score $G_{i}^{n}$ in all tasks.
In this section, we provide a comprehensive visualization of the evolutionary trajectories across all six tasks. Figure 2 plots the round-wise global performance $G^{n}$ for each agent, with dashed lines representing the fitted regression of $S_{\text{evo}}$ to reflect the evolution trend.
  1. 1.
    Universality of Performance Oscillation. The visualization confirms that performance volatility is not task-specific but a universal phenomenon in agent-based evolution. In environments with high strategic complexity, such as Gomoku and Bridge, many agents (e.g., ADK Claude, ADK Gemini) experience significant performance drops in intermediate rounds before potentially recovering. This widespread oscillation reinforces our conclusion in the main text: current agents often fail to reliably distinguish between beneficial and detrimental strategy modifications, leading to the accidental discarding of high-performing code versions.
  2. 2.
    Latent Potential Discovery via Iteration. The trajectories across different tasks suggest that static, zero-shot evaluations often fail to capture an agent’s upper bound. Even in the highly challenging SWE-Perf task, despite the low absolute scores, certain agents exhibit clear upward trajectories, suggesting that iterative feedback allows agents to gradually navigate the complex search space of software engineering.
  3. 3.
    Reliability of the $S_{\text{evo}}$ Metric. The figure highlights the importance of our proposed $S_{\text{evo}}$. While the raw $G^{n}$ is highly sensitive and sometimes suffers from development failures in specific rounds, the dashed regression lines remain relatively stable. This indicates that our evolutionary score provides a more accurate measure of the agent’s evolutionary progress than any single-round performance snapshot.
Appendix H Pairwise Code Similarity Heatmaps
In this section, we provide the detailed pairwise similarity matrices between Round 1 and Round 2 code submissions for all six tasks (Figure 8). In Competitive Tasks, similarity is measured by Action Consistency (%). A value of 100.0 indicates that the agent’s code produces identical behavioral outputs in endgame positions. In objective tasks, similarity is measured by CodeBLEU. These matrices generally show lower absolute values and higher dispersion, reflecting the complexity of syntactic and semantic alignment in large-scale code repositories.
A key observation is that the distribution and range of similarity scores vary significantly by task, reflecting the inherent difficulty of learning and replicating code in different domains. In Hold’em (b), the matrix is characterized by exceptionally high similarity values, including several instances of 100.0% alignment. This indicates that the logic required for this task is relatively "easy to learn" and replicate. Agents can effortlessly adopt a peer’s strategy or maintain their own with perfect consistency, suggesting a lower barrier to entry for strategic imitation in this environment. In Gomoku (a), Bridge (c), and Chess (d), the similarity scores are more dispersed. In objective tasks, even the highest similarity scores rarely exceed 60-70%. This is due to two factors: (1) These tasks involve larger codebases and more intricate dependencies, making it difficult for agents to maintain high syntactic or semantic overlap during refinement. (2) The use of CodeBLEU means that even minor functional improvements or refactoring can lead to a significant drop in measured similarity.